### v3

# T2VUnlearning: A Concept Erasing Method for Text-to-Video Diffusion Models

Abstract
Recent advances in text-to-video (T2V) diffusion models have significantly enhanced the quality of generated videos. However, their capability to produce explicit or harmful content introduces new challenges related to misuse and potential rights violations. To address this newly emerging threat, we propose unlearning-based concept erasing as a solution. First, we adopt negatively-guided velocity prediction fine-tuning and enhance it with prompt augmentation to ensure robustness against prompts refined by large language models (LLMs). Second, to achieve precise unlearning, we incorporate mask-based localization regularization and concept preservation regularization to preserve the model’s ability to generate non-target concepts. Extensive experiments demonstrate that our method effectively erases a specific concept while preserving the model’s generation capability for all other concepts, outperforming existing methods. We provide the unlearned models in https://github.com/VDIGPKU/T2VUnlearning.git.
11footnotetext: Corresponding author.
1 Introduction
With the rapid development of text-to-video (T2V) models [31, 21], it has become possible to generate high-quality videos with a single text prompt. However, as T2V models are often trained on large-scale, unfiltered datasets, they can produce highly realistic yet harmful videos, e.g., inappropriate explicit videos or Deepfake videos of public figures, which may cause serious negative impacts on society. Consequently, preventing the generation of undesirable content by T2V models has become a novel and pressing challenge.
To address this new threat, one might consider retraining T2V models with a filtered dataset. However, this can be extremely expensive and thus infeasible. As an alternative, current approaches rely on prompt manipulation techniques, such as SAFREE [32] which projects toxic tokens in prompts into a harmless subspace. However, SOTA T2V models use LLM-refined prompts rich in details to guide generation by default, which significantly reduces the effectiveness of prompt manipulation.
Figure 1: Illustration of concept erasing results for HunyuanVideo. T2VUnlearning successfully prevent the generation of videos containing the specified target concepts while preserving the knowledge of non-target concepts. Explicit videos (*) have been manually masked for publication.
To alleviate the above issues, we propose T2VUnlearning, a robust and precise concept erasing method for T2V models, which utilizes unlearning [2] to selectively forget knowledge of undesirable concepts. Firstly, we revise the negatively-guided prediction of Text-to-Image (T2I) unlearning methods [12] to accommodate the different generative paradigms of T2V models. Specifically, we fine-tune T2V models with negatively-guided velocity prediction to minimize the probability of generating videos of target concepts, and use prompt augmentation to enhance robustness against LLM refined prompts. Unlike [12] that directly uses the target concept as the generation prompt to generate pseudo-training data, we exploit an uncensored LLaMA3 model [15] to augment the generation prompt with additional context or details, effectively prevent T2V models from generating undesirable contents even when prompted with LLM-refined prompts.
Secondly, we observe that introducing context into the pseudo data can inadvertently cause the model to forget context-related knowledge (e.g., unlearning “dog” with videos of “a dog running on grassy lawn” might also cause the model to forget “lawn”). Hence, we introduce mask-based localization regularization to address this issue. Specifically, we derive attention masks from the text-visual interaction areas of full-attention layers in T2V models [3], and use these masks to localize and constrain unlearning to only affect the target concept like [18].
Furthermore, we observe that unlearning a single concept can cause the model to forget a variety of semantically related non-target concepts (e.g., removing “dog” degrades the model’s ability to generate other animal classes). This phenomenon is similar to Catastrophic Forgetting (CF) [11], where a model forgets previously learned information when learning new knowledge. To prevent the model from forgetting non-target concepts, we propose concept preservation regularization which is inspired by DreamBooth [25]. Specifically, by selecting a semantically related preservation concept, we can significantly enhance the preservation of other non-target concepts, thereby achieving precise unlearning.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to propose an unlearning-based concept erasing method for T2V models.
  • We propose negatively-guided velocity prediction augmented with prompt augmentation, mask-based localization regularization, and concept preservation regularization, to achieve robust and precise concepts erasing for T2V models.
  • Extensive experiments demonstrate that our method significantly outperforms current concept erasing techniques across diverse prompt types and video generation models.
2 Related Work
Text-to-Video Diffusion Models
Following the success of diffusion models in the T2I domain, early T2V works [17, 30, 29, 28, 4] adopted similar architectural paradigms, typically using U-Net backbones with cross-attention mechanisms to inject textual information. For example, VDM [17] employed a 3D U-Net to enhance alignment between text and video features. Building upon this, VideoCrafter2 [5] introduced a two-stage training strategy and decoupled motion and appearance at the data level.
While U-Net-based approaches achieved early success, they struggle with modeling long-range temporal dependencies and suffer from computational inefficiencies. To address these limitations, recent T2V models have shifted toward Diffusion Transformer (DiT) [24] architectures and removed cross-attention. CogVideoX [31] adopts a Multimodal DiT (MMDiT)  [10] that concatenates text and visual tokens as input to a cross-modal 3D full-attention layer, improving both efficiency and performance. HunyuanVideo [21] further explores this direction by introducing dual-stream and single-stream DiT variants: the former processes text and visual tokens separately, while the latter feeds their concatenation into a full-attention layer.
Concept Erasure for Diffusion Models
Diffusion models have achieved remarkable success in both T2I and T2V generation. However, growing concerns over their potential misuse, particularly in producing harmful or NSFW content, have sparked a wave of research on concept erasure, i.e., preventing diffusion models from generating content of specific undesired concepts.
Existing works on concept erasure have focused almost exclusively on the T2I task and can be broadly categorized into two approaches. The first is prompt-manipulation, which leaves the model weights unchanged and instead modifies the input prompt to prevent generating harmful content. For example, SLD [27] introduces a safety guidance term to guide predictions away from target concepts, while SAFREE [32] identifies toxic tokens in the text embedding space and removes them via orthogonal projection, showing promising results on CogVideoX.
The second category is unlearning, which alters the model weights by zero-shot model editing or fine-tuning methods. Zero-shot approaches directly manipulate the model weights—typically within the cross-attention layers—to suppress undesired concepts, as in UCE [13] and RECE [14]. While efficient and easy to apply, these methods rely heavily on cross-attention mechanism, making them incompatible with recent T2V architectures. In contrast, fine-tuning-based methods offer a more generalizable solution. ESD [12] erases concepts by learning to predict noise guided by negative prompts. AdvUnlearn [33] and Receler [18] build upon ESD with adversarial strategies to improve erasure robustness. Despite the progress in the T2I domain, experiments (see Section 4.1) show that existing methods fail to achieve robust and precise concept erasure in the T2V setting.
3 Method
Figure 2: Overview of T2VUnlearning pipeline. T2VUnlearning performs unlearning by inserting adapters after each cross-modal full-attention layer. It generates pseudo data through prompt augmentation, and fine-tunes the model using an unlearning loss $\mathcal{L}_{unlearn}$ to erase the target concept, a mask-based localization regularization $\mathcal{L}_{loc}$, and a preservation regularization $\mathcal{L}_{pre}$ to retain non-target concepts.
3.1 Background
T2V Diffusion Models
Diffusion models [16] learn to generate data by reversing a process that gradually adds Gaussian noise to input. Given an input data point $\mathbf{x}$ and condition $y$, a diffusion model $\epsilon_{\theta}$ is trained to predict the noise $\epsilon$ from a noisy sample $\mathbf{x}_{t}$ at a given time step $t$:
$\mathcal{L}_{\textrm{diffusion}}=\mathbb{E}_{\mathbf{x},y,\epsilon,t}\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{x}_{t},y,t\right)\right\|_{2}^{2}.$ (1)
In T2V diffusion models, the training objectives and generation paradigms often differ. In this work, we focus on two widely used open-source models: CogVideoX [31] and HunyuanVideo [21].
CogVideoX employs v-prediction parameterization [26], where the model is trained to predict a velocity vector. Specifically, given a video latent representation $\mathbf{x}$ and condition $y$, the model adds Gaussian noise $\epsilon$ by $t$ steps and the noisy sample can be expressed as $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon$, where $\alpha_{t}$ and $\sigma_{t}$ are scheduling coefficients. The model $\mathbf{v}_{\theta}$ is encouraged to predict the ground truth velocity vector $\mathbf{v}=\alpha_{t}\mathbf{x}-\sigma_{t}\epsilon$ by the following loss function:
$\mathcal{L}_{\textrm{cogvideox}}=\mathbb{E}_{\mathbf{x},y,\epsilon,t}\left\|\mathbf{v}-\mathbf{v}_{\theta}\left(\mathbf{x}_{t},y,t\right)\right\|_{2}^{2}.$ (2)
HunyuanVideo is based on rectified flow [23], which trains the model to learn a vector field that maps a complex data distribution to a simple Gaussian distribution. Specifically, given a latent video representation $\mathbf{x}$ and condition $y$, it initializes Gaussian noise $\epsilon$ and constructs a training sample $\mathbf{x}_{t}$ at step $t$ by linear interpolation between $\mathbf{x}$ and $\epsilon$. The model $\mathbf{v}_{\theta}$ is trained to predict the velocity $\mathbf{u}_{t}=d\mathbf{x}_{t}/dt$ that guides the noise back to the data by the following loss function:
$\mathcal{L}_{\textrm{hunyuan}}=\mathbb{E}_{\mathbf{x},y,\mathbf{x}_{0},t}\left\|\mathbf{u}_{t}-\mathbf{v}_{\theta}\left(\mathbf{x}_{t},y,t\right)\right\|_{2}^{2},$ (3)
Unlearning in Diffusion Models
For any Diffusion model with weights $\theta$, the goal of unlearning is to reduce the probability of generating video $\mathbf{x}$ described by target concept $c$ [12] by:
$P_{\theta^{{}^{\prime}}}(\mathbf{x})\propto\frac{P_{\theta}(\mathbf{x})}{P_{\theta}(c|\mathbf{x})^{\eta}},$ (4)
where the $\theta^{{}^{\prime}}$ indicate new model weights after unlearning, and $\eta$ is a tunable parameter indicating unlearning strength. Increasing $\eta$ forces the model to lower the probability of generating contents containing target concepts, and the target concepts are more likely to be completely erased.
3.2 Overview of Pipeline
T2VUnlearning is an adapter-based fine-tuning method that achieves robust and precise unlearning through three strategies: negatively-guided velocity prediction with prompt augmentation, mask-based localization regularization, and concept preservation regularization. An overview of the pipeline is illustrated in Figure 2.
3.3 Negatively-guided Velocity Prediction with Prompt Augmentation
Given any T2V model $\mathbf{v}_{\theta}$ with weights $\theta$, our goal is to obtain an unlearned model weight $\theta^{{}^{\prime}}$ that reduces the probability of generating video $\mathbf{x}$ described by target concept $c$. Based on Equation 4, we reformulate the expression in terms of probability density. At diffusion timestep $t$, this leads to:
$p_{\theta^{{}^{\prime}}}({\mathbf{x}_{t}})\propto\frac{p_{\theta}({\mathbf{x}_{t}})}{p_{\theta}(c|{\mathbf{x}_{t}})^{\eta}}.$ (5)
Using Bayes’ theorem $p(c|{\mathbf{x}_{t}})=\frac{p({\mathbf{x}_{t}}|c)p(c)}{p({\mathbf{x}_{t}})}$, we take the logarithm of both sides and compute the gradient with respect to $\mathbf{x}_{t}$}, rewriting it in the form of score prediction:
$\nabla_{\mathbf{x}_{t}}\log p_{\theta^{{}^{\prime}}}({\mathbf{x}_{t}})$ $\propto\nabla_{\mathbf{x}_{t}}\log p_{\theta}({\mathbf{x}_{t}})$ (6)
$-\eta(\nabla_{\mathbf{x}_{t}}\log p_{\theta}({\mathbf{x}_{t}}|c)-\nabla_{\mathbf{x}_{t}}\log p_{\theta}({\mathbf{x}_{t}})),$
which can be reparameterized to the form of velocity prediction. Specifically, for HunyuanVideo with gaussian probability path $p(\mathbf{x}_{t}|\mathbf{x})=N(\mathbf{x}_{t};(1-t)\mathbf{x},t^{2}I)$, there exists an equivalence [22] between velocity $\mathbf{u}_{t}$ and score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$:
$\mathbf{u}_{t}=-\frac{t}{1-t}\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})-\frac{1}{1-t}\mathbf{x}_{t}.$ (7)
This formula implies that knowing $\mathbf{u}_{t}$ allows us to compute $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$, and consequently, the score estimator $\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t})$ can be equivalently converted to the velocity estimator $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$. Similarly, for CogVideoX, given the diffusion noise assumptions and Tweedie’s formula [9], we can show that the score estimator $\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t})$ can be converted to the velocity estimator $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ (see Appendix 2). Therefore, the objective in Equation 6 can be reparameterized to velocity prediction:
$\mathbf{v}_{neg}$ $=\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-\eta(\mathbf{v}_{\theta}(\mathbf{x}_{t},c,t)-\mathbf{v}_{\theta}(\mathbf{x}_{t},t)),$ (8)
$\mathcal{L}_{unlearn}$ $=\mathbb{E}_{\mathbf{x},c,\epsilon,t}\left\|\mathbf{v}_{neg}-\mathbf{v}_{\theta^{{}^{\prime}}}\left(\mathbf{x}_{t},c,t\right)\right\|_{2}^{2},$
where $\mathbf{v}_{neg}$ is the right-hand side of Equation 6 rewritten in terms of velocity. Note that $\mathbf{v}_{neg}$ is similar to negative guidance in classifier-free guidance, thus we refer to it as negatively-guided velocity prediction.
Inspired by ESD [12], we use the original T2V model to generate pseudo data instead of manually collecting real training data. Specifically, $\mathbf{x}_{t}$ at step $t$ is obtained by denoising a sampled Gaussian noise to timestep $t$ instead of adding noise to real data. Previous works generate $\mathbf{x}_{t}$ using only the target concept as the prompt. However, training solely on simple instances of concept $c$ is insufficient to reduce the generation probability for contents with additional context or detail, because these more complex examples come from a different distribution that the model hasn’t seen during fine-tuning.
Figure 3: Example of prompt augmentation.
To address this problem, we propose a simple yet effective solution: augmenting the prompts used to generate $\mathbf{x}_{t}$. As shown in Figure 3, we followed the prompt-rewriting strategies employed during the training of CogVideoX, utilizing an uncensored variant of LLaMA3***https://huggingface.co/Orenguteng/Llama-3-8B-Lexi-Uncensored and apply in-context prompting to generate augmented prompts that align with the format requirements of T2V models. Experimental results (see Section 4.4) demonstrate that this approach achieves robust unlearning when tested against a wide range of paraphrased prompts, including those written by humans.
3.4 Mask-based Localization
Incorporating contextual information into the pseudo data $\mathbf{x}$ can inadvertently cause the model to forget knowledge of those additional contexts and details. Prior work in T2I, such as  Huang et al. [18], addresses this issue by computing a concept-related cross-attention mask and restricting model updates to the masked region. Despite the absence of explicit cross-attention in T2V models, interactions between text and visual tokens still exist. As demonstrated by DiTCtrl [3], the query-key (QK) maps in full-attention layers can be divided into distinct regions, among which the text-to-video region exhibits interactions analogous to cross-attention. Consequently, extracting a concept-related attention mask based on the text-to-video region of the QK map is feasible (as shown in Figure 4).
Figure 4: Example of obtaining attention mask through full-attention QK map.
Therefore, we adapt the mask-based localization technique proposed in Receler [18] to T2V unlearning. Specifically, we employ adapter-based fine-tuning by inserting a linear adapter that modifies visual tokens after each full-attention layer. To achieve localized unlearning, we compute a concept-related attention mask $M$ by thresholding the text-to-video attention map, and constrain the adapter to operate only on the visual tokens within this mask by:
$\mathcal{L}_{loc}=\frac{1}{L}\sum_{l=1}^{L}\left\|o^{l}\odot(1-M)\right\|_{2}^{2},$ (9)
where $L$ denotes the total number of full-attention layers, $o^{l}$ is the output of the adapter at $l$-th layer, and $\odot$ signifies element-wise multiplication.
3.5 Concept Preservation against Catastrophic Forgetting
We observed that erasing a specific concept from T2V models can inadvertently cause the models to forget semantically related concepts. This is particularly pronounced in HunyuanVideo, where unlearning a single concept can negatively impact a wide range of non-target concepts, thereby detrimenting the overall generative capability of the model. This phenomenon is analogous to Catastrophic Forgetting (CF) [11], where a model’s performance on previously learned tasks significantly degrades when fine-tuned on new tasks. T2I fine-tuning method DreamBooth [25] addresses CF by introducing a prior preservation term. Inspired by this, we propose a preservation regularization term to enhance non-target concept preservation in HunyuanVideo. Specifically, for a target concept $c$, we introduce preserve concept $c^{pre}$, and ask the model to preserve generation capability on $c^{pre}$, by minimizing:
$\mathcal{L}_{pre}=\mathbb{E}_{\mathbf{x}^{pre},c^{pre},\epsilon,t}\left\|\mathbf{v}_{\theta}^{{}^{\prime}}\left(\mathbf{x}^{pre}_{t},c^{pre},t\right)-\mathbf{v}_{\theta}\left(\mathbf{x}^{pre}_{t},c^{pre},t\right)\right\|_{2}^{2},$ (10)
where $\mathbf{x^{pre}}$ denotes generated data of preserve concept. We select the concept most likely to be affected by the erasure of the target concept—typically a semantically similar or closely related concept, such as “person” for nudity erasure or a different face for face erasure. Experimental results (see Section 4.4) demonstrate that this preservation regularization substantially improves the generation capabilities of all other non-target concepts.
Combining Equation 89 and 11, our final optimization objective can be formulated as:
$\mathcal{L}=\mathcal{L}_{unlearn}+\alpha\mathcal{L}_{loc}+\beta\mathcal{L}_{pre},$ (11)
where $\alpha$ and $\beta$ is a tunable parameter.
4 Experiment
Figure 5: Comparison on the sexually-explicit subset of SafeSora. We report the total count of frames detected as a certain NudeNet label.
In this section, we demonstrate the effectiveness and robustness of our proposed method through comprehensive experiments. We evaluate T2VUnlearning on three SOTA T2V models: CogVideoX-2B, CogVideoX-5B [31], and HunyuanVideo [21], using publicly available Diffusers-based implementations and pretrained weights. For brevity, we refer to these models as CogX-2B, CogX-5B, and Hunyuan, respectively, in all tables and figures. We mainly compare our method against SAFREEWe adapt SAFREE to HunyuanVideo based on its implementation for CogVideoX, using mean pooling to extract the prompt representation and compute its distance to the toxic subspace. [32], a SOTA prompt manipulation approach, and negative prompting (referred to as NegPrompt), the most widely used technique for discouraging the generation of undesirable contents. Additinally, we adapt unlearning methods of T2I models (i.e., ESD [12] and Receler [18]) for a more comprehensive comparison. Unless otherwise specified, we train each T2V model using the Adam optimizer [20] with a learning rate of $1\times 10^{-4}$ for 500 epochs, and all experiment is conducted with bf16 precision. Additional training details are provided in Appendix 4. We provide the result of erasing nudity in Section 4.1, erasing celebrity faces in Section 4.3, and erasing Imagenet objects in Section 4.2, as well as ablation study in Section 4.4.
4.1 Nudity Erasure
Setting
We evaluate nudity erasure using three datasets: (1) Gen, a set of prompts generated by us that describe nudity with rich context and detail (examples can be seen in Appendix B); (2) Ring-A-Bell, a set of stylized short prompts depicting explicit artwork, adopted from  Yoon et al. [32], Gong et al. [14]; and (3) SafeSora [6], from which we take the subset of human-written sexually explicit prompts. To assess the efficacy of nudity erasure, we generate 49 frames per prompt using each T2V model at its default resolution and apply the NudeNet detector [1] to identify frames containing nudity. We report the Nudity Rate, defined as the proportion of frames labeled with any nudity-related tag by NudeNet.
To analyze the potential impact of nudity erasure on non-nudity concepts, we adopt VBench [19]—a widely used video generation benchmark. We evaluate the nudity-erased model on the Object Class and Subject Consistency metrics to assess its ability to generate videos of benign concepts, as well as its ability to generate temporally coherent videos.
Table 1: Comparison of nudity erasure performance against prior methods. We report the Nudity Rate on the Gen and Ring-A-Bell datasets, along with Object Class and Subject Consistency scores from the VBench benchmark on HunyuanVideo.
[…]
Object
[…]
Subject
[…]
CogX-2B CogX-5B Hunyuan CogX-2B CogX-5B Hunyuan
Original 57.10 61.80 78.08 30.25 42.50 69.85 88.57 95.53
NegPrompt 36.00 46.35 55.35 11.75 14.91 56.73 91.94 93.45
SAFREE 32.43 35.12 48.37 14.23 10.64 50.48 48.48 94.92
Ours 19.73 16.47 12.73 6.97 2.74 20.71 87.00 94.70
Quantitative Results
As shown in Table 1, our method outperforms previous approaches across all prompt sets and T2V models, removing over 76% of nudity content. Its advantage becomes more pronounced when evaluated with longer and more detailed prompts. Furthermore, despite modifying model weights, our method maintains performance comparable with prompt manipulation techniques when generating benign content. It successfully preserves HunyuanVideo’s ability to generate non-nudity objects and ensures strong temporal consistency. In contrast, NegPrompt occasionally introduces overly saturated artifacts, while SAFREE may mistakenly generate human figures in place of non-nudity objects (see Figure 10 in Appendix 5). We further validate our method’s effectiveness on human-written prompts of SafeSora in Figure 5, which presents detailed NudeNet label detection results on videos generated by HunyuanVideo. Notably, our method significantly lowers the probability of generating highly explicit content, such as genitals, buttocks, and female breasts, erasing over 80% of the total detected instances.
Apart from T2V methods, we also compare our approach with unlearning methods for T2I models. As shown in Table 4.1, our method achieves the lowest nudity rate while maintaining comparable or better performance in Object Class, indicating that it effectively erases the target concept without significantly impacting non-target concepts.
Table 2: Comparison agasint T2I methods on nudity erasure. We adapt T2I methods to HunyuanVideo and report Nudity Rate on Gen dataset.
[…]
Original 78.08 88.57
ESD 47.96 79.95
Receler 68.37 89.41
Ours 12.73 87.00
Table 3: Results of face erasure. The Original row reports the ID-Similarity between the output of the original model and the ground-truth identity. The Erase row reports the ID-Similarity of the target face, while the Preserve row shows the average ID-Similarity of the four non-target faces.
Model Methods Merkel Obama Trump Biden Elizabeth AVG
[…]
4.2 ImageNet Object Erasure
Settings
We evaluate object erasure following the protocol of ESD, selecting 10 distinct ImageNet classes [7] as target concepts. Specifically, we erase one concept at a time and evaluate the preservation of the remaining nine. We conduct per-frame classification and calculate average $\text{top}_{\text{k}}$ accuracy, defining the Erasure Success Rate (ESR-k) as $1-\text{top}_{\text{k}}$ accuracy for the erased concept, and the Preservation Success Rate (PSR-k) as the average $\text{top}_{\text{k}}$ accuracy for the preserved concepts.
Experiments are conducted on the CogVideoX-2B model. For each target concept, we fine-tune the model for 300 epochs and generate 17-frame videos using 20 diverse prompts per concept.
Quantitative Results
As demonstrated in Table 4, our method achieves the highest ESR-1 score, demonstrating its ability to prevent the generation of target concepts even under long and detailed prompts. At the same time, it maintains preservation performance comparable to prior prompt-manipulation methods, indicating its effectiveness in retaining knowledge of non-target concepts. Notably, unlike other approaches that fail to achieve a high ESR-5 score, our method achieves a high ESR-5. This indicates that T2VUnlearning is the only method capable of completely erasing or severely distorting the target concept, whereas other methods typically introduce only minor distortions that may cause classification errors but rarely lead to full removal.
To further validate this, we conducted a user study on ImageNet object erasure. To reduce potential bias, we recruited 66 participants with diverse genders, ages, and academic backgrounds. Each participant was shown 15 sets of generated samples from four methods and asked to select the sample in which the target concept was most effectively erased. As illustrated in Figure 6, our method was selected in over 80% of the cases, confirming its superiority in erasing target concepts. We present the full evaluation results as well as the qualitative results in Appendix 6.
Table 4: Results of ImageNet object erasure. We compare the Erasure Success Rate (ESR) and Preservation Success Rate (PSR) on the CogVideoX-2B model.
[…]
Figure 6: User study of ImageNet object erasure.
Figure 7: Visualization of face erasure on HunyuanVideo. Each column corresponds to the output of an unlearned model. The target face for erasure is highlighted with an orange box, while the other faces are to be preserved.
4.3 Face Erasure
Settings
To evaluate a more challenging case of concept erasure and preservation, we apply our method to a face erasure task. Compared to objects, human faces exhibit higher inter-class similarity, increasing the risk that erasing one identity may inadvertently affect others. We select five public figures, setting one identity as the target while preserving the others. We use ID-Similarity as the evaluation metric, calculating the cosine similarity between ArcFace [8] feature embeddings of the generated faces and their corresponding ground-truth identities. For each identity, we generate 17-frame videos using 30 prompts.
Quantitative Results
As demonstrated in Table 3, our method effectively reduces the identity similarity of the target face across multiple T2V models, indicating successful erasure. While erasing faces may inevitably influence the generation of other similar faces, our method minimizes such interference. Notably, HunyuanVideo exhibits strong robustness, where the preservation performance on non-target identities remains largely unaffected. We further presents qualitative results from HunyuanVideo in Figure 7.
4.4 Ablation Study
Prompt Augmentation and Regularization
We apply prompt augmentation to enhance the robustness of unlearning and employ both localization and preservation regularization to achieve precise concept removal. To validate the effectiveness of these strategies, we disable each component individually and evaluate their impact on the nudity-erasure task using HunyuanVideo. As shown in Table 5, removing prompt augmentation significantly degrades the unlearning performance when handling LLM-refined prompts. Removing the regularization terms, on the other hand, negatively impacts non-target concepts. As illustrated in Figure 8, removing localization regularization can enhance nudity erasure, but it comes at the cost of degraded generation quality for non-nudity concepts. Removing preservation regularization has an even more pronounced impact on non-target concepts, leading to degradation across a wide range of benign concepts, as evidenced by the Object Class metric.
Table 5: Ablation results on strategies used to enhance T2VUnlearning. We evaluate the impact of removing Prompt Augmentation (w/o Aug), Localization(w/o $\mathcal{L}_{loc}$), and Preservation (w/o $\mathcal{L}_{pre}$), and report corresponding results.
Nudity
[…]
Object
[…]
Subject
[…]
Ours 12.73 87.00 94.70
w/o Aug 50.69 86.23 94.66
w/o $\mathcal{L}_{loc}$ 0.45 81.84 93.97
w/o $\mathcal{L}_{reg}$ 8.37 68.37 93.47
Unlearning Strength
We perform unlearning by predicting negatively-guided velocity, where the negative guidance scale $\eta$ controls the strength of unlearning, indicating the extent to which the prediction deviates from the original. In practical applications, the unlearning strength $\eta$ can be adjusted based on the specific models and target concepts. In our experiments, we set $\eta=3.0$ for nudity erasure on HunyuanVideo, $\eta=5.0$ for face erasure, and $\eta=7.0$ for all experiments on CogVideoX, and found these settings to be effective. Increasing $\eta$ generally improves erasure efficacy: for example, when unlearning “nudity” in HunyuanVideo, using $\eta=1$ results in a 37.92% Nudity Rate, whereas $\eta=7$ reduces it to 1.22%. However, a higher $\eta$ may also negatively impact the non-target concepts; for instance, setting $\eta=7$ lowers the Object Class metric to 73.52.
Figure 8: Ablation study of different strategies in T2VUnlearning. Top: Generated with prompt “A couple relaxes together, openly nude, in a warm, natural setting”. Bottom: Generated with prompt “a person. Realistic, Natural lighting, Casual”. Images are manually masked for publication.
5 Conclusion
In this paper, we propose T2VUnlearning, the first unlearning-based concept erasure method tailored for T2V models. Our method is robust across diverse target concepts, prompting styles, and model architectures, and achieves precise unlearning by minimizing negative impact on non-target concepts. To effectively eliminate the target concept, we adopt negatively-guided velocity prediction enhanced with prompt augmentation to improve robustness against LLM-refined prompts. Additionally, we introduce mask-based localization regularization and concept preservation regularization to mitigate degradation in the generation of non-target concepts. Extensive experiments validate the effectiveness and generalizability of our method.